b1az3:

The data list shows the cost, in dollars, of different brands of sneakers at a shoe store. 44.99, 59.99, 32.30, 65.00, 52.59 What is the population standard deviation, to the nearest tenth, of the data (assuming this is all of the data)?

AZ:

\(\sigma =\sqrt{ \dfrac{\sum(x_i-\mu)^2}{N}}\) where \(\sigma\) = population standard deviation \(\sum\) = the sum from 1 to N \( x_i\) = each individual data point \(\mu\) = mean of the data list \( N\) = total number of data points so to first find the population standard deviation, can you tell me what the average or mean is? 44.99, 59.99, 32.30, 65.00, 52.59 How many numbers do we have? That's going to be our N. Add them all and divide by that number N. That will be our mean. What do you get?